# ImplicitQA: Going beyond frames towards Implicit Video Reasoning

### v1

@@ H.8 Inferred Counting
H.9 Causal and Motivational Reasoning
Finally, in the Causal and Motivational Reasoning example in Figure 23, GPT-O3 and GPT-4.1 correctly attribute the escape of the rats to being discovered, while others fail to connect the causal event. These examples collectively highlight the diversity and difficulty of implicit reasoning tasks in ImplicitQA.
Figure 21: More Qualitative ImplicitQA examples, targeting distinct implicit-reasoning dimensions.
Figure 22: More Qualitative ImplicitQA examples, targeting distinct implicit-reasoning dimensions.
Figure 23: More Qualitative ImplicitQA examples, targeting distinct implicit-reasoning dimensions.

### v2

@@ H.8 Inferred Counting
H.9 Motivational Reasoning
Finally, in the Motivational Reasoning example in Figure 24, GPT-O3 and GPT-4.1 correctly attribute the escape of the rats to being discovered, while others fail to connect with the relevant event. These examples collectively highlight the diversity and difficulty of implicit reasoning tasks in ImplicitQA.
Figure 22: More Qualitative ImplicitQA examples, targeting distinct implicit‐reasoning dimensions.
Figure 23: More Qualitative ImplicitQA examples, targeting distinct implicit‐reasoning dimensions.
Figure 24: More Qualitative ImplicitQA examples, targeting distinct implicit‐reasoning dimensions.